OpenStudy (anonymous):

Hi guys, please help me to find the possible value of x. x^3+2x^2-29x-30

OpenStudy (lgbasallote):

is there an = 0 at the end there?

OpenStudy (anonymous):

To verify check if f(-1) = 0 \[f(-1) = (-1)^3 + 2(-1)^2 -29(-1) -30 \\ = f(-1) = -1 + 2 +29 -30 = 0\] Hence \[(x+1) \ is \ a \ factor \ of \ x^3 + 2x2 - 29x -30\] < apply remainder theory>

OpenStudy (anonymous):

Hello, DO you know finding a ZERO of a term as such lets take (x+1) What you have to do is - x+1=0 there fore x= -1 ( shifting the terms)

OpenStudy (anonymous):

k thanks yarr